# Towards Robust Automated Perceptual Voice Quality Assessment with Deep Learning

### v1

@@ Abstract
Objective: Perceptual voice quality assessment plays a critical role in diagnosing and monitoring voice disorders by providing standardized evaluation of vocal function. Traditionally, this process relies on expert raters utilizing standard scales, such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and Grade, Roughness, Breathiness, Asthenia, and Strain (GRBAS). However, these metrics are inherently subjective and susceptible to inter-rater variability, which motivates the need for automated and objective assessment methods. Methods: In this study, we propose Voice Quality Assessment Network (VOQANet), a deep learning-based framework with an attention mechanism that leverages a Speech Foundation Model (SFM) to capture high-level acoustic and prosodic information from raw speech for automated voice quality assessment. To further enhance robustness and interpretability, we present VOQANet+, which integrates handcrafted acoustic features such as jitter, shimmer, and harmonics-to-noise ratio (HNR) with SFM embeddings into a hybrid representation. Unlike previous studies that focus only on vowel-based phonation (the PVQD-A subset) on the Perceptual Voice Quality Dataset (PVQD), we evaluate our models on both vowel-based phonation and sentence-level speech (the PVQD-S subset) to improve generalizability in real-world applications. Results: Experimental results show that sentence-based input yields stronger performance than vowel-based input, especially at the patient level, highlighting the value of longer utterances in capturing perceptual voice attributes. VOQANet consistently outperforms the baseline methods in terms of root mean squared error (RMSE) and Pearson correlation coefficient (PCC) across CAPE-V and GRBAS dimensions, while VOQANet+ performs even better. Additional experiments under noisy conditions show that VOQANet+ maintains higher prediction accuracy and robustness. Conclusion: These findings highlight the effectiveness of combining SFM embeddings with domain-informed acoustic features for interpretable and resilient automated voice quality assessment. Significance: VOQANet+ shows strong potential for deployment in real-world and telehealth settings, addressing the limitations of subjective perceptual assessments with an interpretable and noise-resilient solution.

### v2

# Towards Robust Automated Perceptual Voice Quality Assessment with Speech Foundation Models

@@ Abstract
Perceptual voice quality assessment is essential for diagnosing and monitoring voice disorders by providing standardized evaluations of vocal function. Traditionally, expert raters use standard scales such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and Grade, Roughness, Breathiness, Asthenia, and Strain (GRBAS). However, these metrics are subjective and prone to inter-rater variability, motivating the need for automated, objective assessment methods. This study proposes Voice Quality Assessment Network (VOQANet), a deep learning-based framework with an attention mechanism that leverages a Speech Foundation Model (SFM) to extract high-level acoustic and prosodic information from raw speech. To enhance robustness and interpretability, we also introduce VOQANet+, which integrates handcrafted acoustic features—such as jitter, shimmer, and harmonics-to-noise ratio (HNR)—with SFM embeddings into a hybrid representation. Unlike prior studies focused only on vowel-based phonation (PVQD-A subset) of the Perceptual Voice Quality Dataset (PVQD), we evaluate our models on both vowel-based and sentence-level speech (PVQD-S subset) to improve generalizability. Results show that sentence-based input outperforms vowel-based input, especially at the patient level, underscoring the value of longer utterances for capturing perceptual voice attributes. VOQANet consistently surpasses baseline methods in root mean squared error (RMSE) and Pearson correlation coefficient (PCC) across CAPE-V and GRBAS dimensions, with VOQANet+ achieving even better results. Additional experiments under noisy conditions show that VOQANet+ maintains high prediction accuracy and robustness, supporting its potential for real-world and telehealth deployment. These findings demonstrate the effectiveness of combining SFM embeddings with domain-informed acoustic features for interpretable and resilient voice quality assessment.